# SwapMoE: Efficient Memory-Constrained Serving of Large Sparse MoE Models via Dynamic Expert Pruning and Swapping

### v2

@@
Abstract.
Mixture of experts (MoE) is a popular technique to improve capacity of large models with conditionally-activated parallel neural network modules (experts). Due to its remarkable scaling performance with sparse computation, it is widely used in modern Large Language Models (LLMs) and Large Vision Models (LVMs). However, serving such large models on edge devices is challenging due to memory constraints. Typical solutions like memory swapping or weight pruning may lead to significantly higher latency or severe accuracy loss.
In this paper, we introduce SwapMoE, a framework for efficient continuous MoE-based large models serving with tunable memory budgets. The main idea of SwapMoE is to keep a small dynamic set of important experts, namely Virtual Experts, in the main memory for inference, while seamlessly maintaining how the Virtual Experts map to the actual experts. We use a profiling-guided planner to allocate the resources for SwapMoE that can fully utilize the memory budgets and bandwidth, and an importance-aware scheduler to efficiently identify, update, and use the Virtual Experts for accurate inference.
To evaluate SwapMoE, we conduct experiments on multiple edge devices with state-of-the-art MoE-based Large Language Models and Large Vision Models. The results demonstrate remarkable performance of SwapMoE under various memory constraints. Specifically, SwapMoE can enable running large MoE models under tight memory budgets with similar latency to pruned compact models, while with significantly higher accuracy.
submissionid: #399copyright: acmcopyright
1. Introduction
@@
Recently, the world has witnessed the great advancement of deep learning (DL), represented by pre-trained large Transformers. Such an advancement is driven by the phenomenon that the larger model capacity generally leads to higher intelligence (Kaplan et al., 2020). Among various attempts to scale up neural networks, Mixture of Experts (MoE) is a promising technique that can avoid linearly increasing computation based on conditional sparse computation.
An MoE model can be created from a normal model by replacing several layers with MoE layers, each consisting of multiple parallel neural network modules (experts). In each layer, the experts are trained to accommodate different data distributions, and only the most relevant experts are activated (by a gate module) and used for computation for each input sample during the inference process. Due to this benefit, MoE is now widely used in almost all AI applications, including computer vision (Riquelme et al., 2021; Li et al., 2023; Xue et al., 2021), neural language processing (Fedus et al., 2022; Shazeer et al., 2017), and speech recognition (You et al., 2021, 2022).
However, unlike cloud servers that can elastically expand the memory with distributed computing, edge devices often have limited memory, which prevent them to enjoy the benefits of MoE. For example, a Switch Transformer (Fedus et al., 2022) with 64 experts per layer requires 14 GiB of memory for inference, which is impossible to fit in edge devices that typically have only 8 GiB, 4 GiB, or lower memory size. What’s more, many edge devices need to serve multiple applications, and the system-allocated memory budget for each application is even more limited.
@@
Researchers have proposed various methods to reduce the memory consumption of MoE models. The most straightforward approach from the system perspective is to use memory swapping, i.e. dynamically loading/unloading the expert parameters from/to the external memory. For example, Huang et al. (2023) propose to load experts into the memory on demand to reduce memory footprint, EdgeMoE (Yi et al., 2023) focuses on memory swapping in the decoding phase of MoE-based language models and reduces the swapping overhead using quantization. Brainstorm  (Cui et al., 2023) proposed consolidating related experts onto the same device to reduce communication overhead and combining multiple expert kernels into a single kernel for computation, albeit without reducing memory usage. There are also algorithm-perspective approaches (Chen et al., 2022; Kim et al., 2021) that propose to prune the experts to reduce model size. However, swapping-based approaches either have to trade latency for a reduced memory footprint, while pruning-based methods would lead to accuracy loss and require model training. How to efficiently serve an off-the-shelf MoE model under memory constraints remains challenging.
@@
Fortunately, there is a unique property in many MoE workloads that can be exploited for performance optimization - data distribution locality. For example, in generative language models, the output tokens are produced one by one, and they mostly belong to the same semantic context. The successive queries of a user are also semantically related. In object detection, the input images fed into the model are usually continuous video frames. Such locality can potentially lead to consistent activation patterns of MoE experts, therefore create room for more intelligent expert management. Furthermore, the temporal locality property of video analytics has been extensively studied in prior work (Wakatsuki et al., 2021) for accelerating normal CNN models, while how it can be used in MoE for memory reduction is unknown yet.
@@
Achieving efficient and accurate inference under memory constraint with SwapMoE requires to further address two main challenges. The first is to dynamically identify the important experts as the Virtual Experts to use in the main memory. The second is to seamlessly update the Virtual Experts to avoid extra I/O waiting latency and memory consumption.
We adopt a two-phase design, including an online model serving phase and an offline planning phase, and address the aforementioned challenges with various techniques in both phases. In the online phase, the job of SwapMoE is to handle the whole inference process with the small set of Virtual Experts, update the Virtual Experts to keep the most important experts in memory, and make sure the update process is non-blocking. The main difficulty to maintain the Virtual Experts is to infer the importance of each expert (including the off-memory ones) without using the expert for actual computation. We find a lightweight indicator based on the understanding of MoE operations. During inference, we modify the gating logic to redirect the control flows from the original experts to the Virtual Experts, and update the Virtual Experts with lightweight, amortized, asynchronous parameter loading.
The goal of the offline phase is to plan the execution strategy of SwapMoE (including the per-layer memory allocation and expert update frequency) to meet the memory constraints. The total number of experts in Virtual Experts is determined by the memory budget, but how to divide the budget to different layers is based on each layer’s tolerance of absent experts. We empirically find that different models may have different patterns of tolerance, which can be modeled with a neural network. After determining the memory allocation, we further calculate the update frequency based on the profiled I/O bandwidth - we always try to maximize the update frequency as long as the I/O operations does not block the computation.
We implement SwapMoE in two popular frameworks for large MoE models, including Huggingface Transformers (Wolf et al., 2019) and Tutel (Hwang et al., 2022). To evaluate the effectiveness of SwapMoE, we conduct experiments with Swin-MoE (Liu et al., 2022), Switch Transformer (Fedus et al., 2022) (SwitchT for short), and GPTSAN (gpt, 2023) models on both computer vision and natural language processing tasks, compared with the normal model inference scheme and strong baselines. For example, on text summarization tasks with Switch Transformer, SwapMoE can reduce the memory consumption from 14.2 GiB to 4.7 GiB (67% less), together with 50% latency reduction and a slight Rouge-2 score drop of 0.041.
These results have shown that SwapMoE enables the serving of large MoE models on resource-constrained edge devices with limited memory budgets.
@@
  • We propose a novel MoE inference framework that supports efficient continuous serving of large MoE models under memory constraints. Our method enables deployment of off-the-shelf MoE models to edge devices with tunable memory.
@@
  • We implement SwapMoE with popular inference frameworks and conduct extensive experiments with large MoE models on edge devices. The results have demonstrated the effectiveness of our approach on popular computer vision and neural language processing tasks.
2. Background and Motivation
2.1. MoE Preliminaries
Figure 1. Illustrations of MoE layers. MoE layers are usually used with Transformer architectures where the layer input is a sequence of tokens.
Increasing the capacity of a model can improve its performance (Kaplan et al., 2020). A prevailing method for scaling up models is using MoE, where certain model layers are replaced as parallel experts and are sparsely activated for inference based on their scores given by the gating network. In GPT-MoE (Artetxe et al., 2022), it is observed that some experts within the model exhibit a higher proficiency in English, while others demonstrate a greater aptitude for handling French inputs.
A typical MoE module is shown in Figure 1(b). Let $G(x)$ denote the output of the gating network for a given input $x$, and let $E_{i}(x)$ denote the output of the $i$-th expert network for the same input $x$. Then, the output $y$ of the MoE module can be expressed as:
(1) $y=\sum_{i=1}^{n}G(x)_{i}E_{i}(x),$
where $n$ is the number of expert networks in the MoE module, and $G(x)_{i}$ is the gating weight assigned to the $i$-th expert network for the input $x$. Most existing MoE models are based on the transformer structure (Vaswani et al., 2017). Therefore, the input $x$ in Equation 1 represents a token, and each layer of the MoE model receives multiple tokens as input. The gating weights $G(x)_{i}$ are typically calculated using a softmax function over a set of learned parameters. We can convert any layer of a deep learning model into an MoE module.
@@ 2.1. MoE Preliminaries
Among different MoE variants, the sparse MoE (Shazeer et al., 2017) is the most commonly used, in which only one or few (represented as $k$) experts are activated for each input. Specifically, in a sparse MoE with $k=1$, only the $i$-th expert with max $G(x)_{i}$ is activated for input $x$, and the output $y=G(x)_{i}E_{i}(x)$.
@@
The MoE structure is often accompanied by the Transformer architecture (Liang et al., 2022; Zhang et al., 2022), in which the input of each MoE layer is a sequence of tokens and each token may choose different experts in one MoE layer, as shown in Figure 1(b).
2.2. Memory Constraints of Edge AI
In the past decade, the growth in memory capacity of mobile/edge devices has significantly lagged behind the advances in computing power. For example, from iPhone 4 in 2010 to iPhone 14 in 2022, the GPU and CPU performance of iPhone series has increased by nearly 100-1000 times (in terms of FLOPS) (John, 2016), but the RAM capacity has only increased by a factor of 8 (GadgetVersus, 2023).
Unlike cloud servers that can elastically expand the memory with distributed interconnected GPUs, the memory capacity of edge devices is usually a hard constraint. The AI models that exceed the memory constraints are usually infeasible to run on edge devices. For instance, the memory size of most edge devices today typically range from 1-8 GiB, while the model sizes of many state-of-the-art MoE-based language models and vision models are larger than 8 GiB.
Meanwhile, the memory constraints not only come from the hardware, but also from the operating systems and applications. Many edge devices have to serve multiple applications and each commercial application has multiple internal AI-based functions, so the memory budget for each AI model is further limited.
Figure 2. The memory footprint of each layer in SwitchT-32 (i.e. Switch Transformer with 32 experts per MoE layer). 89.42% of the memory is consumed by the MoE layers (even layers).
Therefore, reducing the memory footprint of MoE models is desirable to fit them into memory-constrained edge devices. As shown in Figure 2, most of the memory is consumed by the MoE layers, which is the bottleneck to optimize.
2.3. Problems of Conventional Solutions
@@
Existing methods for reducing the resource overhead of MoE model inference include on-demand loading (Lane et al., 2016) (i.e. loading the parameters of MoE layers when they are needed and release afterwards) and MoE model pruning (Chen et al., 2022; Kudugunta et al., 2021) (i.e. cutting the less important experts permanently).
@@
While the on demand loading method can reduce the GPU memory usage during MoE model inference without affecting model accuracy, it introduces significant latency overhead with each MoE layer parameter transmission via PCIe. In Figure 3, running MoE models with on-demand loading introduces 6.2x-8.9x higher latency. Meanwhile, the transfer rate of PCIe is approximately 10 GiB/s, whereas the large MoE model’s parameters are at least 1 GiB. Our measured IO bandwidth usage is 10.2 GiB/s, while the latency for one model inference is approximately 0.2 s. As a result, the transmission of a large amount of parameter data can significantly impede the computation process. As shown in Figure 4 and Figure 5, model parameters transmission takes up most inference time and significantly obstructs the model’s computational process and may deplete IO resources.
@@
Figure 3. Naive on-demand expert loading reduces memory, but also results in huge inference overhead.
Figure 4. Latency breakdown of MoE model inference with layer-wise memory swapping. The transmission of model weights consumes the majority of the time.
Figure 5. Weight loading may block computation when running MoE with layer-wise memory swapping. Due to the large size of MoE layers and sparse computation, loading the weights of a layer is always slower than computing the layer, which slows down the inference even if the weights are loaded asynchronously.
2.4. Data Distribution Locality in MoE Serving
@@
Data distribution locality is an important characteristic in many edge AI applications, which refers to the phenomenon that successive input samples are distributionally similar or correlated. The locality property is common in both language modeling and video analytic tasks that MoE models are mainly used for. Firstly, language modeling tasks are naturally subject to locality - each inference process of the model produces a token, and the successive tokens belong to the same sentence. Secondly, edge AI models are usually deployed in a fixed environment and used for serving an individual user or organization, the successive input samples (e.g. video frames, conversations) are semantically close to each other.
@@
Since the experts in MoE models are designed to handle different data distributions, there exists an opportunity to cache the most relevant experts in the main memory at each time step, therefore reducing the memory consumption. The data distribution locality in edge applications further produces the chance to reuse the cached experts for successive input samples, which can reduce the overhead of parameter loading.
Such locality properties have been frequently and suceesfully exploited for edge AI inference acceleration (Wakatsuki et al., 2021; Kong et al., 2023). For example, one can skip or simplify the inference process of some video frames based on previous inference results in video analytic tasks. These approaches inspired our design, but are orthogonal to our work. In this paper, we aim to exploit the locality to optimize memory-constrained serving of MoE models, including large language models. Both the memory aspect and MoE models are rarely discussed in prior locality-based optimizations.
3. Our Design: SwapMoE
Figure 6. The workflow of SwapMoE. Given an off-the-shelf MoE model and a memory-constrained edge device, we satisfy the constraint by executing the model with a smaller set of experts (Virtual Experts). The Virtual Experts are selected, used and updated seamlessly at runtime, and the memory allocation for the experts is determined at offline.
We introduce SwapMoE to achieve efficient and accurate MoE inference under tunable memory budgets. The main idea is to perform MoE inference with a subset of “Virtual Experts” in memory (which reduces the memory footprint) and seamlessly update the Virtual Experts according to their relevance to current data distribution (which maintains accuracy and latency). The key technical challenges to be addressed in SwapMoE include:
  1. (1)
    How to select and use the Virtual Experts in an off-the-shelf MoE model, so that the MoE model can produce accurate prediction with a small set of Virtual Experts.
  2. (2)
    How to dynamically manage the Virtual Experts via seamless memory operations that can minimize latency/memory overhead at runtime.
3.1. Overview
SwapMoE addresses the challenges with various techniques in a two-phase holistic design, as shown in Figure 6.
In the online phase, the job of SwapMoE is to efficiently identify, update, and use a subset of experts (Virtual Experts) for memory-constrained MoE inference. We utilize a lightweight magnitude-based method to estimate the importance score of each expert, and maintain the most important experts in memory (i.e. Virtual Experts). These experts are used to handle the whole inference process, which will be discussed in Section 3.2. Furthermore, SwapMoE needs to dynamically updates the Virtual Experts to maintain important experts seamlessly (without impacting the latency), which will be discussed in Section 3.3.
The job of the offline phase is to obtain an optimal memory plan for the online phase to facilitate efficient and accurate MoE inference. To achieve this, we first obtain several important performance metrics and models through fine-grained expert profiling, including memory footprint, compute latency, loading latency, and model accuracy, which will be discussed in Section 3.4. The profiling needs to be conducted only once for a specific device. Then the profiling-guided expert planner utilizes the established relations to derive the optimal memory plan (layer space allocation and expert I/O frequency) for Virtual Experts based on a given memory budget, which will be discussed in Section 3.5.
3.2. Importantance-aware Virtual Experts Selection & Inference
@@
To maintain accuracy and latency with Virtual Experts-based MoE inference, we must wisely select the most important experts and update them according to the data distribution to maintain accuracy with Virtual Experts-based MoE inference. To ensure that the Virtual Experts remain up-to-date and consist of the most important experts for the current data distribution, the first question is how to determine the importance of each expert at runtime. To address this problem, we introduce the concept of the expert importance score.
The expert importance score is an input-dependent metric, denoted as $importance(E_{i},X)$, that approximates the positive contribution of the expert $E_{i}$ to the inference process of the input $X$. Keeping the experts with higher importance scores in the Virtual Experts are expected to produce more precise predictions with the MoE model.
[…]
Based on the mechanism of MoE in Equation 1, we notice that each expert $E_{i}$ influences the output of the MoE layer by adding its output $G(x)_{i}E_{i}(x)$ to the output. Naturally, we speculate that the key factor determining the importance of $E_{i}$ is the value of $||G(x)_{i}E_{i}(x)||$, where |||||$\cdot$||| | ⋅ | | represents the magnitude. However, we cannot directly calculate $||G(x)_{i}E_{i}(x)||$ for each expert during the inference process, because the non-Virtual Experts experts are not available in the main memory to do the inference of $E_{i}(x)$. As a result, we need to approximate the value of $||G(x)_{i}E_{i}(x)||$ without calculating $E_{i}(x)$.
After analyzing Equation 1, we approximate $||G(x)_{i}E_{i}(x)||$ using $||x||*||G(x)_{i}||*||E_{i}||$ to determine the importance of experts, where $||E_{i}||$ is the parameter magnitude of $E_{i}$ that can be pre-computed offline. The rationale behind this approximation is that the larger $||x||$, $||G(x)_{i}||$ and $||E_{i}||$ typically lead to higher $||G(x)_{i}E_{i}(x)||$ in high-dimention matrix multiplication, and vice versa. Therefore, the final formula to calculate the importance of experts is
(2) $importance(E_{i},X)=\sum_{x}^{\mathcal{X}_{i}}||x||*||G(x)_{i}||*||E_{i}||,$
@@
Next, we explain how SwapMoE perform inference only with the selected Virtual Experts in main memory. Once Virtual Experts has been selected, we can calculate the output of the MoE layer, as defined by Equation 1.
Figure 7. (a) Original gating of MoE: all experts may by used for inference; (b) Ours Masked Gating: only Virtual Experts will be used.
As shown in Figure 7, we use Masked Gating to redirect all inference requests to Virtual Experts. Specifically, once the computation of $G(x)$ in Equation 1 is completed, we achieve the flow of all tokens’ inference requests to Virtual Experts by assigning scores corresponding to unimportant experts as $-inf$:
$G(x)_{i}=-inf,\text{if i not in Virtual Experts},$
then the function $G(x)$ will be adjusted through the softmax operation. In essence, by performing the max operation on $softmax(G(x))$, none of the unimportant experts will be selected.
3.3. Seamless Virtual Experts Update
To run Virtual Experts-based MoE inference efficiently, we need to seamlessly perform the expert updates without latency overhead. Once we have the importance scores for all experts, we can update Virtual Experts accordingly at runtime by loading the important experts into the main memory and the unimportant experts out. In our design, the number of experts in the Virtual Experts is fixed, so we only need to load the parameters of new experts to overwrite the old experts.
The most critical question that arises in expert loading is the additional I/O cost, which may cause idle waiting time before computation and significantly slow down the inference process. We introduce two techniques to reduce this overhead, including amortized expert loading and asynchronous expert loading, as shown in Figure 8 (ii).
Amortized expert loading.
Figure 8. Different loading strategy of Virtual Experts. i) synchronously updating all experts after single sample inference; ii) amortized asynchronous expert loading across different samples.
According to Equation 2, the importance scores of all experts in all MoE layers may be updated for each input sample. One nature way is loading experts for all layers and every sample during inference. This method results in optimal accuracy, but the overall system latency is unacceptably high due to the large number of experts in the MoE model. For instance, in the SwitchT-16 model (Switch Transformer with 16 experts per MoE layer), if we update 50% of the experts, approximately 5GB of data needs to be transferred (typical data transfer speed of GPUs is approximately 10 GiB/s), resulting in high latency. Therefore, we adopt a fixed frequency for updating, which is determined by the update frequency set in the profiling phase.
However, simply reducing the frequency of updates is not enough. If we do the loading of expert periodically, the samples that undergo expert loading would suffer from high inference latency due to the extensive data transmission and limited bandwidth between the main memory and external memory (or disk). While other samples that do not undergo expert loading will have lower inference latency. Thus, we further distribute this periodic overhead of expert parameter transmission across different input samples to fully utilize the I/O bandwidth. Specifically, in each update cycle that spans $k$ samples, we evenly partition the pending experts (that are determined to load into Virtual Experts) to $k$ samples. Within every $\frac{k}{\text{\# of moe layers}}$ we just load experts of one MoE layer. In this way, we can distribute the cost of expert loading to samples.
Asynchronous expert loading. The amortized loading strategy can reduce the size of per-sample data transmission, but it will still increase the inference latency due to the additional I/O waiting time. Fortunately, we observe that replacing experts in Virtual Experts does not affect the inference computation process. Therefore, we adopt a computing while transferring strategy, where the actual replacement of the new experts is performed asynchronously during the inference computation of old experts. Once the loading process of new experts is done, the inference process is notified to use the new experts and expire the old experts. While the I/O cost of expert loading is not changed in this asynchronous design, the delay in the overall inference latency is mitigated by the concurrent nature of transferring and computing, ensuring the efficient execution of the model.
3.4. Fine-grained Expert Profiling

### v3

# SwapMoE: Serving Off-the-shelf MoE-based Language Models with Tunable Memory Budget

@@
Abstract
Mixture of experts (MoE) is a popular technique to improve capacity of Large Language Models (LLMs) with conditionally-activated parallel experts. However, serving MoE models on memory-constrained devices is challenging due to the large parameter size. Typical solutions such as memory swapping or expert pruning may lead to significantly higher latency or severe accuracy loss. In this paper, we introduce SwapMoE, a framework for efficient serving of MoE-based large language models with tunable memory budgets. The main idea of SwapMoE is to keep a small dynamic set of important experts, namely Virtual Experts, in the main memory for inference, while seamlessly maintaining how the Virtual Experts map to the actual experts. Experiments have shown that SwapMoE can reduce the memory footprint while maintaining reasonable accuracy. For example, on text summarization tasks with Switch Transformer, SwapMoE can reduce the memory consumption from 14.2 GiB to 4.7 GiB, together with 50% latency reduction and a slight Rouge-2 score drop of 0.041.
SwapMoE: Serving Off-the-shelf MoE-based Language Models with Tunable Memory Budget
Rui Kong1,2thanks:  Work was done while the author was interning at Institute for AI Industry Research (AIR), Tsinghua University.  Yuanchun Li2,3thanks:  Corresponding authors: Yuanchun Li, Linghe Kong.  Qingtian Feng2,411footnotemark: 1  Weijun Wang2 Xiaozhou Ye5  Ye Ouyang5  Linghe Kong122footnotemark: 2  Yunxin Liu2,3 1Shanghai Jiao Tong University 2Institute for AI Industry Research (AIR), Tsinghua University 3Shanghai Artificial Intelligence Laboratory 4National University of Singapore 5AsiaInfo Technologies (China), Inc.
1 Introduction
@@
Recently, the world has witnessed the great advancement of pre-trained large language models Li et al. (2024). Such an advancement is driven by the phenomenon that the larger model capacity generally leads to higher intelligence Kaplan et al. (2020). Among various attempts to scale up neural networks, Mixture of Experts (MoE) Shazeer et al. (2017) is a promising technique that can avoid linearly increasing computation based on conditional sparse computation. However, memory-costrainted devices is often a major concern in edge AI training and serving Huang et al. (2023b); Gim and Ko (2022); Wang et al. (2022); Ma et al. (2023); Kong et al. (2023). The constraint is even more challenging for large MoE models. For example, a Switch Transformer Fedus et al. (2022) with 64 experts per layer requires 14 GiB of memory for inference, which is impossible to fit in consumer devices that typically have only 8 GiB, 4 GiB, or lower memory size. What’s more, many consumer devices need to serve multiple applications, and the system-allocated memory budget for each application is even more limited.
@@
Researchers have proposed various methods to reduce the memory footprint of MoE model inference. The most straightforward approach is to use memory swapping, i.e. dynamically loading/unloading the parameters from/to the external memory. For example, Huang et al. (2023a) propose to load experts into the memory on demand to reduce memory footprint but introduce additional latency overhead, EdgeMoE Yi et al. (2023) focuses on memory swapping in the decoding phase of MoE-based language models and reduces the swapping overhead using quantization. However, it only considers the scenario of single token inference, which does not align with real-world use cases. Other algorithm-perspective approaches Chen et al. (2022); Kim et al. (2021) propose to prune the experts to reduce model size. However, swapping-based approaches either have to trade latency for a reduced memory footprint, while pruning-based methods would lead to accuracy loss and require model training. How to efficiently serve an off-the-shelf MoE model under memory constraints remains challenging.
@@
Fortunately, there is a unique property in many MoE workloads that can be exploited for performance optimization - activation locality. For example, in generative language models, the output tokens are produced one by one, and they mostly belong to the same semantic context. The successive queries of a user are also semantically related. Such locality can potentially lead to consistent activation patterns of MoE experts, therefore create room for more intelligent expert management.
@@
We implement SwapMoE with Huggingface Transformers Wolf et al. (2019) library. To evaluate the effectiveness of SwapMoE, we conduct experiments with Switch Transformer Fedus et al. (2022) (SwitchT for short), and GPTSAN gpt (2023) models on natural language processing tasks, compared with the normal model inference scheme and strong baselines. For example, on text summarization tasks with Switch Transformer, SwapMoE can reduce the memory consumption from 14.2 GiB to 4.7 GiB (67% less), together with 50% latency reduction and a slight Rouge-2 score drop of 0.041, which shows that SwapMoE enables the serving of large MoE models on resource-constrained consumer devices with limited memory budgets.
@@
  • We propose a novel MoE inference framework that supports efficient serving of large MoE models under memory constraints. Our method enables deployment of off-the-shelf MoE models to consumer devices with tunable memory budgets.
@@
  • We implement SwapMoE with popular inference frameworks and conduct extensive experiments with large MoE models on consumer devices. The results have demonstrated the effectiveness of our approach on natural language processing tasks.
2 Background and Motivation
@@
The sparse MoE Shazeer et al. (2017) is the most commonly used, in which only one or few (represented as $k$) experts are activated for each input. Specifically, in a sparse MoE with $k=1$, only the $i$-th expert with max $G(x)_{i}$ is activated for input $x$, and the output $y=G(x)_{i}E_{i}(x)$.
@@
The MoE structure is often accompanied by the Transformer architecture Liang et al. (2022); Zhang et al. (2022), in which the input of each MoE layer is a sequence of tokens and each token may choose different experts in one MoE layer.
2.1 Limitations of Conventional Solutions
Figure 1: SwitchT-16: Naive on-demand expert loading reduces memory, but also results in huge inference overhead.
Figure 2: SwitchT-16: Latency breakdown of MoE model inference with layer-wise memory swapping. The transmission of model weights consumes the majority of the time.
Figure 3: Weight loading may block computation when running MoE with layer-wise memory swapping. Due to the large size of MoE layers and sparse computation, loading the weights of a layer is always slower than computing the layer, which slows down the inference even if the weights are loaded asynchronously.
Figure 4: The workflow of SwapMoE. Given an off-the-shelf MoE model and a memory-constrained consumer device, we satisfy the constraint by executing the model with a smaller set of experts (Virtual Experts). The Virtual Experts are selected, used and updated seamlessly at runtime, and the memory allocation for the experts is determined at offline.
@@ 2.1 Limitations of Conventional Solutions
Existing methods for reducing the resource overhead of MoE model inference include on-demand loading Lane et al. (2016) (i.e. loading the parameters of MoE layers when they are needed and release afterwards) and MoE model pruning Chen et al. (2022); Kudugunta et al. (2021) (i.e. cutting the less important experts permanently).
@@
While the on demand loading method can reduce the GPU memory usage during MoE model inference without affecting model accuracy, it introduces significant latency overhead with each MoE layer parameter transmission via PCIe. In Figure 1, running MoE models with on-demand loading introduces 6.2x-8.9x higher latency. Meanwhile, the transmission of a large amount of parameter data can significantly impede the computation process. As shown in Figure 2 and Figure 3, model parameters transmission takes up most inference time and significantly obstructs the model’s computational process and may deplete IO resources.
@@
2.2 Activation Locality in MoE Models
@@
Data distribution locality is an important characteristic in many AI applications, which refers to the phenomenon that successive input samples are distributionally similar or correlated. Firstly, each inference process of the model produces a token, and the successive tokens belong to the same sentence. Secondly, AI models are usually deployed in a fixed environment and used for serving an individual user or organization, the successive input samples (e.g. conversations) are semantically close to each other.
@@
Since each expert in a MoE model is trained to handle certain data distribution, there exists an opportunity to cache the most relevant experts in the main memory at each time step, therefore reducing the memory consumption. The data distribution locality in AL applications further produces the change to reuse the cached experts for successive input samples, which can reduce the overhead of parameter loading.
3 Our Design: SwapMoE
Figure 5: (a) Original gating of MoE: all experts may by used for inference; (b) Ours Masked Gating: only Virtual Experts will be used.
Figure 6: Different loading strategy of Virtual Experts. i) synchronously updating all experts after single sample inference; ii) amortized asynchronous expert loading across different samples.
SwapMoE utilizes a two-phase holistic design, as shown in Figure 4. In the online phase, the job of SwapMoE is to efficiently identify, update, and use a subset of experts (Virtual Experts) for memory-constrained MoE inference. The job of the offline phase is to obtain an optimal memory plan for the online phase to facilitate efficient and accurate MoE inference.
3.1 Importantance-aware Virtual Experts Selection & Inference
@@
Firstly, we must wisely select the most important experts and update them according to the data distribution to maintain accuracy with Virtual Experts-based MoE inference. To address this problem, we introduce the concept of the expert importance score, which can be written as
$importance(E_{i},X)=\sum_{x}^{\mathcal{X}_{i}}||x||*||G(x)_{i}||*||E_{i}||,$ (1)
@@
As shown in Figure 5, we use Masked Gating to redirect all inference requests to Virtual Experts.
3.2 Seamless Virtual Experts Update
Once we have the importance scores for all experts, we can update Virtual Experts accordingly at runtime by loading the important experts into the main memory and the unimportant experts out. We introduce two techniques to reduce overhead, including amortized expert loading and asynchronous expert loading, as shown in Figure 6 (ii).
3.3 Fine-grained Expert Profiling